OpenStudy (anonymous):

Write an equation of a line that passes through the point (3, 2) and is parallel to the line y = 3x - 4.

OpenStudy (anonymous):

) y = 3x + 7 2) y = 3x - 7 3) y = x + 2 4) y = x - 2

OpenStudy (triciaal):

parallel lines have the same slope given line y = 3 x - 4 the slope is ? this is the slope of the parallel line the line passes through (3, 2) and (x, y) equation of line in point-slope format (y-2) = m(x-3) m = slope from above

OpenStudy (anonymous):

wouldn't it be y=3x-3

OpenStudy (triciaal):

y = mx + b using (3,2) and m = 3 2 = 3(3) + b b = -7 y = 3x -7 option 2

OpenStudy (anonymous):

ohh I see now, thanks so much

OpenStudy (triciaal):

welcome